Logiweb diagnose of opgave

Up Help

No errors found

The pyk compiler, version 0.grue.20060417+ by Klaus Grue GRD-2006-06-28.UTC:01:17:25.716427 = MJD-53914.TAI:01:17:58.716427 = LGT-4658174278716427e-6